I have the following question Shown below. I really need help with it please. I have tried working it out but don't know if its right. I tried multiplying 8.3 x 10^13 square feet and 200 feet. then divide by 1.3 x 10^-9 but my answer does not show in my choices for the question. Any help will be greatly appreciated.

The sahara desert covers about 8.3 x 10^13 square feet.
The average depth of the sand is 200 ft
A grain of sand has a volume of approxemately 1.3 x 10^-9
best estimate of the number of grains of sand in the sahara desert?
Thank you

Answer by scott8148(6628)

(8.3 x 10^13)(200) = 1660 x 10^13 = 1.66 x 10^16

(1.66 x 10^16) / (1.3 x 10^-9) = 1.28 x 10^25 (approx)


FYI: there may be an issue with significant figures
___ 200 (no decimal point) has one significant figure
___ "technically" , the answer should have only one significant figure

also , what are the dimensions for the sand grain volume? ___ there may be a units conversion issue